The acid dissociation constant Ka of alloxanic acid HC4H3N2O5 is ×2.2410−7. Calculate the pH of a 0.57M solution of alloxanic acid. Round your answer to 1 decimal place.
Answer #1

Answer

pH = 3.4

Explanation

Ionization equilibrium of alloxanic acid is

HC4H3N2O5(aq) + H2O(l) <-------> C4H3N2O5-(aq) + H2O(l)

Ka = [C43N2O5-] [ H3O+] / [HC4H3N2O5] = 2.24 ×10-7

Initial concentration

[HC4H3N25] = 0.57

[C4H3N2O5-] = 0

[H3O+] = 0

change in concentration

[HC4H3N2O5] = -x

[C4H3N2O5-] = + x

[H3O+] = +x

so,

x2/( 0.57 - x) = 2.24 ×10-7

x = 0.0003572

[H3O+] = 0.0003572M

pH = -log( 0.0003572M)

pH = 3.4